HCl(aq) + Mg(OH)2(aq) = MgCl2(aq) + H2O(l)
1
Expert's answer
2020-10-14T05:47:49-0400

Given reaction is,

HCl(aq) + Mg(OH)2(aq) = MgCl2(aq) + H2O(l)


It's a unbalanced equation,

So, we need to balance them

In a balanced chemical reaction number of each atoms in the reactants is equals to the number of each atoms in the products.



Number of Cl atoms in the both sides are unbalanced , we can balance them by putting '2' in the stoichiometric of the HCl.

And balance the other compound accordingly.


The chemical equation will be,

2HCl(aq) + Mg(OH)2(aq) = MgCl2(aq) + 2H2O(l)


Hence the balance chemical equation is,

2HCl(aq) + Mg(OH)2(aq) = MgCl2(aq) + 2H2O(l)